# Understanding How Synthetic Impurities Affect Glyphosate Solubility and Crystal Growth Using Free Energy Calculations and Molecular Dynamics Simulations

### v1

@@ Introduction
For the DC simulations, we place a bulk of crystal glyphosate pre-equilibrated on one side of the simulation box, and water (or a glyphosate solution) on the remaining space of the simulation box avoiding intermolecular overlapping between both phases frenkel2023understanding ; espinosa2013fluid ; sanchez2024predictions ; sanchez2021parasitic ; garaizar2024toward . We carry out two types of DC simulations: (i) including a perfect crystal of glyphosate in contact with a glyphosate solution; (ii) a glyphosate crystal with random vacancies introduced at the interface in contact with the solution. Once the system reaches equilibrium—normally after 100 nanoseconds—we compute density profiles along the long axis of DC simulation box using the GROMACS tool gmx density gromacs . Then the solubility ($m_{glyphosate}$) is estimated as:
@@
Glyphosate has been provided by the Bayer Crop Science department, with HPLC grade  95% purity, and was further recrystallized to yield HPLC grade  99% purity. Glycine and Phosphoric Acid have been purchased from Sigma-Aldrich Chemie GmbH, Germany, with HPLC grade  99% purity, while deionized water has been obtained directly at our laboratory using a Milli-Q iQ system. We measure the solubility and detection time using a Crystal16 (Technobis Crystallization Systems, 2021). Here the detection time is defined as the time elapsed between reaching the supersaturation point and the first detection of crystals. Solubilities are measured according to the following protocol: in a vial selected amounts of glyphosate, glycine, phosphoric acid, and water ware weighed and allowed to equilibrate at 293 K for three hours. The temperature is then gradually increased at a rate of 1 K/min up to 333 K, and the point at which the solution becomes clear (the clear point) was interpreted as the saturation temperature. For the water-glyphosate system, the solubility measurements obtained via the clear point method were also confirmed by equilibrating a suspension at 304 K for 24 hours, and measuring the concentration of the liquid phase via HPLC.
@@
$t_{\text{growth},j}=\min_{i\in I_{j}}\left[t_{ji}\right],$
@@ I.1 Simulations capture crystal density but underestimate glyphosate solubility
Initially, we assess the force field’s ability to reproduce well-established properties of glyphosate, such as the crystal density, by preparing a $4\times 4\times 4$ replication of the crystal unit cell knuuttila1979crystal (Fig. 1a, left) at 300 K and 1 bar in the $NpT$ ensemble. We then study how the crystal density fluctuates throughout a simulation and compare it to the experimental data obtained by Wilson et al. wilson2023discerning , obtaining values in reasonable agreement (Fig. 1a, right panel). Moreover, we determine that the crystal structure is well conserved throughout the MD simulation, preserving the crystal’s structural integrity.
Once we confirm that the crystal is accurately modelled and that correctly maintains its crystal structure, we continue by preparing DC initial configurations, where the crystal structure is in direct contact with different aqueous solutions, which contain varying concentrations of glyphosate in the solution. We corroborate that the initial glyphosate concentration in the solution does not affect the obtained solubility determination by performing multiple simulations with different initial concentrations of glyphosate. This is because in the NpT ensemble, the system spontaneously evolves towards equilibrium conditions, in which the crystal coexists with the solution at the saturation concentration espinosa2016calculation . The prepared DC configurations contained 8000 water molecules and the necessary varying glyphosate replicas in the solution ranging from 0 mol/kg to 0.690 mol/kg.
@@
Furthermore, we perform DC simulations to examine how the presence of vacancies affect the kinetics to reach the equilibrium concentration of glyphosate in the solution. To build these systems, we randomly remove 10 glyphosate molecules from the outer crystal plane in contact with the solution, artificially creating defects on the surface. The same initial concentrations of the glyphosate solution depicted in Fig. 1c are then used to determine the solubility from the density profiles once the systems reach equilibrium (i.e., the concentration of glyphosate in the solution shows no drift as a function of time). The obtained average results from these two types of simulations using multiple trajectories with different initial concentrations are summarized in Table 1 (further details on these calculations are shown in Figures S1, S2 and Table S1). Based on these results, we do not find significant differences neither in the kinetics nor in the equilibrium value of the DC simulations in presence vs. absence of interfacial defects. Nevertheless, we note that in most of our defect-free simulations, the glyphosate molecules incorporated from solution into the crystal do not adopt the correct bulk lattice structure.
Table 1: Comparison of the experimental wilson2023discerning and computational predictions for the crystal density and solubility limit (using DC simulations with a crystal phase in absence versus presence of vacancies at the interface).
@@ I.2 Calculation of glyphosate solubility using DC simulations exposing different crystal faces
As previously discussed, glyphosate crystallizes from aqueous solution in a pH-dependent zwitterionic form, adopting a monoclinic unit cell. This structure belongs to the $P2_{1}/c1$ space group in the Hermann–Mauguin notation. As shown in Fig. 2a, crystal faces with different Miller indices exhibit distinct characteristics, such as the number of molecules directly exposed to the solvent or their interfacial free energy with the solution. In Section A, we have focused on systems where the [010] crystal face of glyphosate was in contact with the aqueous phase. In the present section, we investigate whether exposing a different crystal face, as illustrated in Fig. 2a, may affect the determination of solubility. From a thermodynamic standpoint, the solubility limit of a crystalline solid in solution is independent of which crystal face is exposed. At equilibrium, the chemical potentials of the solid and dissolved species are equal, and solubility is therefore a bulk property of the coexisting phases rather than a surface-dependent quantity frenkel ; espinosa2016calculation . Nevertheless, different faces present distinct interfacial free energies and growth rates montero2019ice , which can give rise to anisotropic growth and dissolution rates Freitas2020 ; sanchez2021parasitic . In practice, such differences may influence kinetics, even though they should not alter the equilibrium solubility itself. Moreover, comparing solubility values obtained from simulations using different crystal orientations provides an important internal consistency check. If calculations with distinct faces converge to the same solubility, this rules out the possibility that finite-size effects or kinetic trapping are biasing the results, since it is highly unlikely that two independent orientations would be affected by the same artifact in an identical manner. Thus, studying multiple orientations not only probes potential kinetic effects but also strengthens confidence in the robustness of the solubility determination.
@@
To investigate potential kinetic trapping and finite-size effects in our systems, we consider an additional crystal orientation, [001], and construct two new systems for DC simulations. In both cases, the number of water molecules was fixed at 8000, and the required number of glyphosate molecules in the solution was added to achieve initial concentrations of $0\>\text{m}$ and $0.69\>\text{m}$. As shown in Fig. 2b, systems differing in both the exposed crystal plane and the initial glyphosate concentration converge to the same solubility value, indicated by the dashed line. Statistical analysis of the average solubilities for both orientations (Fig. 2c) shows that the results are equivalent within the associated uncertainties. In summary, the crystal face in contact with the aqueous solution does not affect the solubility of glyphosate within the uncertainty, consistent with previous observations in other systems using DC simulations espinosa2016calculation . This agreement between orientations indicates that our simulations are not kinetically trapped and that finite-size effects are negligible.
@@ I.3 Glycine impurities increase glyphosate solubility
Real systems are more complex than just water and glyphosate, as they often contain glycine, a ubiquitous impurity introduced during glyphosate synthesis. At the pH where glyphosate crystallizes, glycine exists mainly in two forms: as a zwitterionic species and a negatively charged species, in a 1:4 ratio yu2002glycine . To examine the effect of this impurity, we construct four systems with varying glycine concentrations while maintaining a fixed number of water and glyphosate molecules in the solution. In addition, we consider both crystal orientations previously described in Sections A and B of the Results. As before, we analyze the DC simulations and obtain the density profiles of the new systems, shown in Fig. 3a. Glycine dynamically adsorbs onto the crystal surface through a continuous adsorption–desorption process, as evidenced by the two maxima highlighted in the inset of Fig. 3a. This behavior is observed for both glycine forms (zwitterionic and negatively charged; see Fig. S4). The snapshot in Fig. 3a further reveals that, instead of aggregating into a solid-like structure on the glyphosate crystal, glycine forms a partial coating layer at the interface, characteristic of transient adsorption. In Fig. 3b, we display the time-evolution of glyphosate concentration converging to its solubility value in the presence (2 wt% glycine) and absence of glycine. The presence of glycine increases the solubility by approximately an order of magnitude. Importantly, this enhancement in solubility cannot be explained solely by the transient surface coating observed in Fig. 3a hindering glyphosate attachment into the surface. We hypothesize that glycine adsorption may result from its concentration exceeding the solubility limit, leading to heterogeneous nucleation at the glyphosate–water interface. However, a detailed investigation of glycine solubility is beyond the scope of this study.
@@
Therefore, the presence of glycine makes it more favorable for glyphosate to remain in solution compared to the impurity-free system. Moreover, this analysis reveals an inverse correlation between glycine concentration and solvation free energy: as glycine concentration increases, the solvation free energy decreases. This result explains the increase in solubility observed in Fig. 3b when glycine is added. In this case, the presence of impurities directly alters the chemical potential of glyphosate in solution. The change in solvation free energy is therefore linked to the solubility of the system through Eq. 2. This mechanism is illustrated in Fig. 4b, where we compare the different systems studied. We observe that the presence of glycine increases glyphosate solubility by approximately an order of magnitude. Comparison of our simulations with experimental data suggests that glycine, frequently present in glyphosate samples zhou2012study , may significantly influence solubility measurements. Using Eq. 2 and the 2 wt% glycine data, we estimate the solvation free energy corresponding to the experimentally reported solubility ($62.4\pm 0.1$ mmol/kg). Assuming a linear correlation between glycine concentration and solvation free energy (Fig. 4a), we predict an effective glycine content of $3.7\pm 0.7$ wt% in the experimental system assuming that the OpenFF would predict the correct solubility for glyphosate. Although the actual glycine content of the experimental samples was not reported gly_sol , this estimation provides a plausible explanation for the observed discrepancy in solubility—beyond force field deficiencies—and further strengthens the value of our computational framework for characterizing and predicting glyphosate phase behavior.
@@ I.4 Experimental validation of the simulation predictions
Figure 4: a) Bar plot of the obtained solubility in the different studied systems. The experimental value is shown in red gly_sol . Simulations without impurities for different crystal planes are shown in blue, whereas simulations where glycine is present as an impurity at different concentrations are shown in green. b) Solvation free energy as a function of glycine weight percentage from free energy calculations. We observe that increasing glycine content leads to a further decrease in solvation free energy.

### v2

@@ ABSTRACT
Glyphosate, crystallization, molecular dynamics simulations, solubility, glycine
[…]
@@ Introduction
For the DC simulations, we place a bulk of crystal glyphosate pre-equilibrated on one side of the simulation box, and water (or a glyphosate solution) on the remaining space of the simulation box avoiding intermolecular overlapping between both phases frenkel2023understanding ; espinosa2013fluid ; sanchez2024predictions ; sanchez2021parasitic ; garaizar2024toward . The exposed crystal orientation is either the (010) or the (001), which we specify in each corresponding section. We carry out two types of DC simulations: (i) including a perfect crystal of glyphosate in contact with a glyphosate solution; (ii) a glyphosate crystal with random vacancies introduced at the interface in contact with the solution. Once the system reaches equilibrium—normally after 100 nanoseconds—we compute density profiles along the long axis of DC simulation box using the GROMACS tool gmx density gromacs . Then the solubility ($m_{glyphosate}$) is estimated as:
@@
Glyphosate has been provided by the Bayer Crop Science department, with HPLC grade  95% purity. We measure the solubility and detection time using a Crystal16 (Technobis Crystallization Systems, 2021). Here the detection time is defined as the time elapsed between reaching the supersaturation point and the first detection of crystals. Solubilities are measured according to the following protocol: in a vial selected amounts of glyphosate, glycine, phosphoric acid, and water ware weighed and allowed to equilibrate at 293 K for three hours. The temperature is then gradually increased at a rate of 1 K/min up to 333 K, and the point at which the solution becomes clear (the clear point) was interpreted as the saturation temperature. For the water-glyphosate system, the solubility measurements obtained via the clear point method were also confirmed by equilibrating a suspension at 304 K for 24 hours, and measuring the concentration of the liquid phase via HPLC.
@@
$t_{\text{growth},j}=\min_{i\in I_{j}}\left[t_{ji}\right],$ (3)
@@ I.1 Simulations capture crystal density but underestimate glyphosate solubility
Initially, we assess the force field’s ability to reproduce well-established properties of glyphosate, such as the crystal density, by preparing a $4\times 4\times 4$ replication of the crystal unit cell knuuttila1979crystal (Fig. 1a, left) at 300 K and 1 bar in the $NpT$ ensemble. We then study how the bulk crystal density fluctuates throughout a simulation and compare it to the experimental data obtained by Wilson et al. wilson2023discerning , obtaining values in reasonable agreement (Fig. 1a, right panel). The small overestimation of the simulated density ($\sim$2%) is within the typical accuracy of a general-purpose fixed-charge force field boothroyd2023development , which has not been explicitly parametrized for a single molecule (e.g., glyphosate) crystal properties. Moreover, we determine that the crystal structure is well conserved throughout the MD simulation, preserving the crystal’s structural integrity and density over time.
Once we confirm that the crystal is accurately modelled and that correctly maintains its crystal structure, we continue by preparing DC initial configurations exposing the (010) crystal orientation, where the crystal structure is in direct contact with different aqueous solutions, which contain varying concentrations of glyphosate in the solution. We corroborate that the initial glyphosate concentration in the solution does not affect the obtained solubility determination by performing multiple simulations with different initial concentrations of glyphosate. This is because in the NpT ensemble, the system spontaneously evolves towards equilibrium conditions, in which the crystal coexists with the solution at the saturation concentration espinosa2016calculation . The prepared DC configurations contained 8000 water molecules and the necessary varying glyphosate replicas in the solution ranging from 0 mol/kg to 0.690 mol/kg (Further details of system sizes can be found in SM section SXIX).
@@
Furthermore, we perform DC simulations, also exposing the (010) orientation, to examine how the presence of vacancies affect the kinetics to reach the equilibrium concentration of glyphosate in the solution. To build these systems, we randomly remove 10 glyphosate molecules from the outer crystal plane in contact with the solution, artificially creating defects on the surface. The same initial concentrations of the glyphosate solution depicted in Fig. 1c are then used to determine the solubility from the density profiles once the systems reach equilibrium (i.e., the concentration of glyphosate in the solution shows no drift as a function of time). The obtained average results from these two types of simulations using multiple trajectories with different initial concentrations are summarized in Table 1 (further details on these calculations are shown in Figures S1, S2 and Table S1). Based on these results, we do not find significant differences neither in the kinetics nor in the equilibrium value of the DC simulations in presence vs. absence of interfacial defects. Nevertheless, we note that in most of our defect-free simulations, the glyphosate molecules incorporated from solution into the crystal do not adopt the correct bulk lattice structure.
Table 1: Comparison of the experimental densitywilson2023discerning and solubilitygly_sol predictions, along with computational calculations (using DC simulations with a crystal phase in absence versus presence of vacancies at the interface).
@@ I.2 Calculation of glyphosate solubility using DC simulations exposing different crystal faces
As previously discussed, glyphosate crystallizes from aqueous solution in a pH-dependent zwitterionic form, adopting a monoclinic unit cell. This structure belongs to the $P2_{1}/c1$ space group in the Hermann–Mauguin notation. As shown in Fig. 2a, crystal faces with different Miller indices exhibit distinct characteristics, such as the number of molecules directly exposed to the solvent or their interfacial free energy with the solution. In Section A, we have focused on systems where the (010) crystal face of glyphosate was in contact with the aqueous phase. In the present section, we investigate whether exposing a different crystal face, as illustrated in Fig. 2a, may affect the determination of solubility. From a thermodynamic standpoint, the solubility limit of a crystalline solid in solution is independent of which crystal face is exposed. At equilibrium, the chemical potentials of the solid and dissolved species are equal, and solubility is therefore a bulk property of the coexisting phases rather than a surface-dependent quantity frenkel ; espinosa2016calculation . Nevertheless, different faces present distinct interfacial free energies and growth rates montero2019ice , which can give rise to anisotropic growth and dissolution rates Freitas2020 ; sanchez2021parasitic . In practice, such differences may influence kinetics, even though they should not alter the equilibrium solubility itself. Moreover, comparing solubility values obtained from simulations using different crystal orientations provides an important internal consistency check. If calculations with distinct faces converge to the same solubility, this rules out the possibility that finite-size effects or kinetic trapping are biasing the results, since it is highly unlikely that two independent orientations would be affected by the same artifact in an identical manner. Thus, studying multiple orientations not only probes potential kinetic effects but also strengthens confidence in the robustness of the solubility determination.
@@
To investigate potential kinetic trapping and finite-size effects in our systems, we consider an additional crystal orientation, (001), and construct two new systems for DC simulations. In both cases, the number of water molecules was fixed at 8000, and the required number of glyphosate molecules in the solution was added to achieve initial concentrations of $0\>\text{m}$ and $0.69\>\text{m}$. As shown in Fig. 2b, systems differing in both the exposed crystal plane and the initial glyphosate concentration converge to the same solubility value, indicated by the dashed line. Statistical analysis of the average solubilities for both orientations (Fig. 2c) shows that the results are equivalent within the associated uncertainties. In summary, the crystal face in contact with the aqueous solution does not affect the solubility of glyphosate within the uncertainty, consistent with previous observations in other systems using DC simulations espinosa2016calculation . This agreement between orientations indicates that our simulations are not kinetically trapped and that finite-size effects are negligible.
@@ I.3 Glycine impurities increase glyphosate solubility
Real systems are more complex than just water and glyphosate, as they often contain glycine, a ubiquitous impurity introduced during glyphosate synthesis. At the pH where glyphosate crystallizes, glycine exists mainly in two forms: as a zwitterionic species and a negatively charged species, in a 1:4 ratio yu2002glycine . To examine the effect of this impurity, we construct four systems with varying glycine concentrations while maintaining a fixed number of water and glyphosate molecules in the solution. In addition, we consider both crystal orientations previously described in Sections A and B of the Results. As before, we analyze the DC simulations and obtain the density profiles of the new systems, shown in Fig. 3a. Glycine is dynamically adsorbed onto the crystal surface through a continuous adsorption–desorption process, as evidenced by the two maxima highlighted in the inset of Fig. 3a. This behavior is observed for both glycine forms (zwitterionic and negatively charged; see Fig. S4). The snapshot in Fig. 3a further reveals that, instead of aggregating into a solid-like structure on the glyphosate crystal, glycine forms a partial coating layer at the interface, characteristic of transient adsorption. We further characterize this behaviour by measuring the surface concentration $\Gamma=N_{\mathrm{Glycine}}/A$ for the (010) and (001) planes at two glycine concentrations (0.5 and 2 wt%). Our surface concentration measurements through DC simulations are summarized in Table 2. Based on our calculations, for a given concentration, the different exposed crystal planes exhibit similar adsorption capacities, as the values of $\Gamma$ agree within the associated uncertainties. Nevertheless, higher glycine concentration leads to a greater surface concentration, a behaviour that is known to further hinder crystal growth in similar systems salvalaglio2012uncovering ; tanaka2022crystal . This prediction will be later confirmed in results from Section D, where we experimentally determine the crystal growth rate as a function of different glycine concentrations.
Plane $\Gamma$ (molecules/nm2)
0.5 % wt Glycine 2 % wt Glycine
(010) 0.11(10) 0.36(10)
(001) 0.18(10) 0.43(10)
Table 2: Surface concentration of glycine ($\Gamma$) for two different crystal orientations of glyphosate measured at two glycine concentrations using DC simulations.
In Fig. 3b, we display the time-evolution of glyphosate concentration converging to its solubility value in the presence (2 wt% glycine) and absence of glycine. The presence of glycine increases the solubility by approximately an order of magnitude. Importantly, this enhancement in solubility cannot be explained solely by the transient surface coating observed in Fig. 3a hindering glyphosate attachment into the surface. We hypothesize that glycine adsorption may result from its concentration exceeding the solubility limit, leading to heterogeneous nucleation at the glyphosate–water interface. However, a detailed investigation of glycine solubility is beyond the scope of this study.
@@
Figure 4: a) Solvation free energy as a function of glycine weight percentage from free energy calculations. We observe that increasing glycine content leads to a further decrease in solvation free energy. b) Bar plot of the obtained solubility in the different studied systems, with different glycine concentrations. The horizontal dashed line depicts the experimental solubility value reported in absence of impurities gly_sol .
Therefore, the presence of glycine makes it more favorable for glyphosate to remain in solution compared to the impurity-free system. Moreover, this analysis reveals an inverse correlation between glycine concentration and solvation free energy: as glycine concentration increases, the solvation free energy decreases. This result explains the increase in solubility observed in Fig. 3b when glycine is added. In this case, the presence of impurities directly alters the chemical potential of glyphosate in solution. The change in solvation free energy is therefore linked to the solubility of the system through Eq. 2. This mechanism is also corroborated in Fig. 4b, where we compare the solubility obtained with increasing glycine concentrations. We observe that the presence of glycine increases glyphosate solubility monotonically by approximately an order of magnitude within the studied concentration range. Moreover, we also observe that, as expected, the solubility is not affected by the crystal orientation exposed to the solution, as already established in Results Section B. Additionally, from our DC simulations, we computed the radial distribution functions between the different heavy atoms of glyphosate and glycine in order to determine which specific interactions are responsible of their intermolecular association that drives the solvation of glyphosate and adsorption of glycine to the crystal. These distributions are shown in Figure S7, and reveal that electrostatic interactions are crucial, being that between glyphosate’s oxygen and glycine’s nitrogen the most relevant one increasing both the solvation free energy (in absolute value) and the solubility as glycine concentration increases (see SM Figure S7).
Further comparison of our simulations with experimental data suggests that glycine, frequently present in glyphosate samples zhou2012study , may significantly influence solubility measurements. Using Eq. 2 and the 2 wt% glycine data, we estimate the solvation free energy corresponding to the experimentally reported solubility ($62.4\pm 0.1$ mmol/kg). Assuming a linear correlation between glycine concentration and solvation free energy (Fig. 4a), we predict an effective glycine content of $2.7\pm 0.5$ wt% in the experimental system assuming that the OpenFF would predict the correct solubility for glyphosate. Although the actual glycine content of the experimental samples was not reported gly_sol , this estimation provides a plausible explanation for the observed discrepancy in solubility—beyond force field deficiencies—and further strengthens the value of our computational framework for characterizing and predicting glyphosate phase behavior.